​Parallelogram ABCD​ is a rectangle.

AC=6y−1
BD=4y+13

What is the value of y?

Answer by ikleyn(52855)

.

The diagonals of any rectangle are congruent,  therefore


    6y - 1  = 4y + 13

    6y - 4y = 13 + 1

      2y    = 14

       y    = 14/2 = 7.    ANSWER

Solved.